Luis punts a football straight up into the air. it is in the air a total of 5 seconds. How many seconds did it spend going up? coming down? and how fast did it leave his foot? How fast was it moving right before hitting the ground, and how high did it go?
1
Expert's answer
2010-09-16T04:11:35-0400
Due to the symmetry of the motion it spent going up 2.5 sec, coming down - 2.5 sec. Use the formula v-gt=0, where t is the time spent going up or down, thus v=gt=10*2.5=25 m/sec. The velocity right before hitting the ground is the same. The height of motion would be:
h=vt-gt^2/2=25*2.5-10*2.5^2/2=31.25 m